You get in line at a popular tobacco shop and notice that there are many people in line ahead of you. During the next 10 minutes you notice that 10 more customers have arrived. You now count 18 people ahead of you in line. Use Little's Law to estimate the total amount of time (in minutes) you should expect to wait to be served.
Answer #1

Flow rate, R = 10/10 = 1 customer per minute

Number of people in line, L = 18

Using little's law, L = W*R ,  where W is expected waiting time

W = L / R

= 18/1

= 18 minutes

Total amount of time expected to wait, W = 18 minutes